A pilot-scale disc-track centrifuge is tested for recovery of bacteria. The centrifuge contains 25 discs with inner and outer diameters 2cm and 10cm, respectively. The half-cone angle is 35 degree. When operated at a speed of 3000 rpm with a feed rate of 3.5 litres/min, 70% of the cells are recovered. If a bigger centrifuge is to be used for industrial treatment of 80 litres/min, what operating speed is required to achieve the same sedimentation performance if the larger centrifuge contains 55 discs with outer diameter 15cm, inner diameter 4.7cm, and half-cone angle 45 degrees?


1
Expert's answer
2021-01-18T01:38:12-0500

 For small scale centrifuge:

D1 = 0.02m , R1 = 0.01m

D0 = 0.10m , R0 = 0.05m

n1 = 25discs

ω1 = 3000rpm

Q1 = 3.5 L/min

θ1 = 35⁰

Large scale centrifuge:

D0,2 = 0.15m , R0,2 = 0.075m

D1,2 = 0.047m , R1,2 = 0.0235m

n2 = 55discs

ω2 = ?

Q2 = 80 L/min

θ2 = 45⁰


Q = "v"g  Σ 

Since the dimensions of particles does not change in both cases , therefore "v"g remains constant.

Q ∝Σ  "\implies" "\frac{Q_1}{ Σ_1}" = "\frac{Q_2}{ Σ_2}"

 Σ1 = "\frac{2\pi nω}{3g}" (R"_0^3" - R"_1^3" ) cotθ


Σ 1 = "\frac{2\pi(25)(2\pi\frac{rad}{rev}3000\frac{rev}{min}\frac{1min}{60s})^2}{3(9.81m/s²)}" (0.053 --0.013 )m3 × "\frac{1}{tan35⁰}"


Σ1 = 106.6m²


Σ2 = "\frac{Q_2}{Q_1}" Σ1

= "\frac{80}{35}"× 106.6

Σ2 = 2436.5m²

Next we find ω2

Σ2 = "\frac{2\pi n_2(\omega_2)^2}{3g}" ( R"_{0,3} ^3" - R"_{1,2}^3" ) cot45⁰


"\omega"2 = "\sqrt{\frac{3gΣ_2}{2\pi n_2(R_{0,1}^3-R_{1,2}^3)cot45⁰}}"

"\omega"2 = "\sqrt{\frac{3(9.81m/s²)2436.5m²}{2
\pi(55)(0.075³-0.0235³)m³cot45⁰}}"

"\omega"2 = 658 "\frac{rad}{s}" × "\frac{rev}{2\pi rad}" × "\frac{60s}{1min}"

"\omega"2 = 6286.7 rpm